# Learning under Imitative Strategic Behavior with Unforeseeable Outcomes

### v1

@@ Abstract
Machine learning systems have been widely used to make decisions about individuals who may best respond and behave strategically to receive favorable outcomes, e.g., they may genuinely improve the true labels or manipulate observable features directly to game the system without changing labels. Although both behaviors have been studied (often as two separate problems) in the literature, most works assume individuals can (i) perfectly foresee the outcomes of their behaviors when they best respond; (ii) change their features arbitrarily as long as it is affordable, and the costs they need to pay are deterministic functions of feature changes. In this paper, we consider a different setting and focus on imitative strategic behaviors with unforeseeable outcomes, i.e., individuals manipulate/improve by imitating the features of those with positive labels, but the induced feature changes are unforeseeable. We first propose a Stackelberg game to model the interplay between individuals and the decision-maker, under which we examine how the decision-maker’s ability to anticipate individual behavior affects its objective function and the individual’s best response. We show that the objective difference between the two can be decomposed into three interpretable terms, with each representing the decision-maker’s preference for a certain behavior. By exploring the roles of each term, we further illustrate how a decision-maker with adjusted preferences can simultaneously disincentivize manipulation, incentivize improvement, and promote fairness.
@@ 1 Introduction
Individuals subject to algorithmic decisions often adapt their behaviors strategically to the decision rule to receive a desirable outcome. As machine learning is increasingly used to make decisions about humans, there has been a growing interest to develop learning methods that explicitly consider the strategic behavior of human agents. A line of research known as strategic classification studies this problem, in which individuals can modify their features at costs to receive favorable predictions. Depending on whether such feature changes are to improve the actual labels genuinely (i.e., improvement) or to game the algorithms maliciously (i.e., manipulation), existing works have largely focused on learning classifiers robust against manipulation (Hardt et al., 2016a) or designing incentive mechanisms to encourage improvement (Kleinberg and Raghavan, 2020; Bechavod et al., 2022). A few studies (Miller et al., 2020; Shavit et al., 2020; Horowitz and Rosenfeld, 2023) also consider the presence of both manipulation and improvement, where they exploit the causal structures of features and use structural causal models to capture the impacts of feature changes on labels.
To model the interplay between individuals and decision-maker, most existing works adopt (or extend based on) a Stackelberg game proposed by Hardt et al. (2016a), i.e., the decision-maker publishes its policy, following which individuals best respond to determine the modified feature. However, these models (implicitly) rely on the following two assumptions that could make them unsuitable for certain applications: (i) individuals can perfectly foresee the outcomes of their behaviors when they best respond; (ii) individuals can change their features arbitrarily at costs, which are modeled as deterministic functions of the feature.
@@
Consider an example of college admission, where the students’ exam scores are treated as features in admission decisions. To get admitted, students may increase their scores by either cheating on exams (manipulation) or working hard (improvement). Here (i) individuals do not know the exact values of their original features (unrealized scores) and the modified features (actual score received in an exam) when they best respond, but they have a good idea of what those score distributions would be like from their past experience; (ii) the cost of manipulation/improvement is not a function of feature change (e.g., students may cheat by hiring an imposter to take the exam and the cost of such behavior is more or less fixed). As the original feature was never realized, we cannot compute the feature change precisely and measure the cost based on it. Therefore, the existing models do not fit for these applications.
@@
Motivated by the above (more examples are also given in App. B.2), this paper studies strategic classification with unforeseeable outcomes. We first propose a novel Stackelberg game to model the interactions between individuals and the decision-maker. Compared to most existing models (Jagadeesan et al., 2021; Levanon and Rosenfeld, 2022), ours is a probabilistic framework that models the outcomes and costs of strategic behavior as random variables. Indeed, this framework is inspired by the models proposed in Zhang et al. (2022); Liu et al. (2020), which only considers either manipulation (Zhang et al., 2022) or improvement (Liu et al., 2020); our model significantly extends their works by considering both behaviors. More importantly, we focus on imitative strategic behavior where individuals manipulate/improve by imitating the features of those with positive labels, due to the following:
@@
  • It is inspired by imitative learning behavior in social learning, whereby new behaviors are acquired by copying social models’ behavior. It has been well-supported by literature in psychology and social science (Bandura, 1962, 1978). Recent works (Heidari et al., 2019; Raab and Liu, 2021) in ML also model individuals’ behaviors as imitating/replicating the profiles of their social models to study the impacts of fairness interventions.
@@
  • Decision-makers can detect easy-to-manipulate features (Bechavod et al., 2021) and discard them when making decisions, so individuals can barely manipulate their features by themselves without changing labels. A better option for them is to mimic others’ profiles. Such imitation-based manipulative behavior is very common in the real world (e.g., cheating, identity theft) and even becomes increasingly worrying during recent years [^1] As COVID-19 hit the world, candidates are more commonly permitted to take exams/assessments (e.g., GRE, TOEFL, or online assessments of companies) remotely. Although many institutions are diligent in designing novel challenges to prevent candidates from directly finding the answers on the internet, the remote nature makes it easier to hire qualified imposters to take the assessments instead of them. Talha (2024) illustrated how students can let others take the GRE instead of them when the test is permitted to be taken at home..
@@
    We study how adjusting the decision-maker’s preferences can affect the optimal policy and its fairness property, as well as the resulting individual’s best response (Sec. 4).
@@ 2 Problem Formulation
Consider a group of individuals subject to some ML decisions. Each individual has an observable feature $X\in\mathbb{R}$ and a hidden label $Y\in\{0,1\}$ indicating its qualification state (“0" being unqualified and “1" being qualified). [^2] Similar to prior work (Zhang et al., 2022; Liu et al., 2018), we present our model in one-dimensional feature space. Note that our model and results are applicable to high dimensional space, in which individuals imitate and change all features as a whole based on the joint conditional distribution $P_{X|Y}$ regardless of the dimension of $X$. The costs can be regarded as the sum of an individual’s effort to change features in all dimensions. Let $\alpha:=\Pr(Y=1)$ be the population’s qualification rate, and $P_{X|Y}(x|1)$, $P_{X|Y}(x|0)$ be the feature distributions of qualified and unqualified individuals, respectively. A decision-maker makes binary decisions $D\in\{0,1\}$ (“0" being reject and “1" being accept) about individuals based on a threshold policy with acceptance threshold $\theta\in\mathbb{R}$: $\pi(x)=P_{D|X}(1|x)=\textbf{1}(x\geq\theta)$. To receive positive decisions, individuals with information of policy $\pi$ may behave strategically by either manipulating their features or improving the actual qualifications. [^3] We assume individuals have budgets to either manipulate or improve. The generalization of considering the actions of “manipulate”, “improve”, and “do nothing” is discussed in App. B.3. Formally, let $M\in\{0,1\}$ denote individual’s action, with $M=1$ being manipulation and $M=0$ being improvement.
Outcomes of strategic behavior. Both manipulation and improvement result in the shifts of feature distribution. Specifically, for individuals who choose to manipulate, we assume they manipulate by “stealing" the features of those qualified (Zhang et al., 2022), e.g., students cheat on exams by hiring qualified imposters. Moreover, we assume the decision-maker can identify the manipulation behavior with probability $\epsilon\in[0,1]$ (Estornell et al., 2021). Individuals, once getting caught manipulating, will be rejected directly. For those who decide to improve, they work hard to imitate the features of those qualified (Bandura, 1962; Raab and Liu, 2021; Heidari et al., 2019). With probability $q\in[0,1]$, they improve the label successfully (overall $\alpha$ increases) and the features conform the distribution $P_{X|Y}(x|1)$; with probability $1-q$, they slightly improve the features but fail to change the labels, and the improved features conform a new distribution $P^{I}(x)$. Throughout the paper, we make the following assumption on feature distributions.
@@
Assumption 2.1 is relatively mild and has been widely used (e.g., (Tsirtsis et al., 2019; Zhang et al., 2020b)). It can be satisfied by a wide range of distributions (e.g., exponential, Gaussian) and the real data (e.g., FICO data used in Sec. 5). It implies that an individual is more likely to be qualified as feature value increases. Meanwhile, compared to the unqualified individuals, the individuals who improve but fail also tend to have higher feature values. Individuals have a good knowledge of their true qualifications by observing their peers or previous individuals who received positive decisions Raab and Liu (2021), and only unqualified individuals have incentives to take action Dong et al. (2018) since $P_{X|Y}(x|1)$ is always the best attainable outcome (as manipulation and improvement only bring additional cost but no benefit to qualified individuals).
@@
The above formulation captures the imitative strategic behavior with unforeseeable outcomes (e.g., college admission example in Sec. 1): individuals best respond based on feature distributions but not the realizations, and the imitation costs (e.g., hiring an imposter) for individuals from the same group follow the same distribution (Liu et al., 2020), as opposed to being a function of feature changes.  (1) above can further be written based on CDF of $C_{M}-C_{I}$, i.e., the difference between manipulation and improvement costs. We make the following assumption on its PDF.
@@
Under Assumption 2.2, $P_{M}(\theta)$ is continuous and satisfies the following: (i) If $q+\epsilon\geq 1$, then $P_{M}(\theta)$ strictly increases. (ii) If $q+\epsilon<1$, then $P_{M}(\theta)$ first increases and then decreases, thereby existing a unique maximizer $\theta_{max}$. Moreover, the maximizer $\theta_{max}$ increases in $q$ and $\epsilon$.
@@ 2.2 Decision-maker’s optimal policy
Strategic optimal policy. Given cost and feature distributions, a strategic decision-maker can anticipate an individual’s best response ((1)) and incorporate it in determining its optimal policy. Under a threshold policy $\pi(x)=\textbf{1}(x\geq\theta)$, the objective $U(\pi)$ can be written as a function of $\theta$, i.e.,
@@ 3 Decomposition of the Objective Difference
As shown in  (4), the objective difference $\Phi$ can be decomposed into three terms $\phi_{1},\phi_{2},\phi_{3}$. It turns out that each term is interpretable and indicates the impact of a certain type of individual behavior on the decision-maker’s utility. We discuss these in detail as follows.
@@
    Benefit from the successful improvement ${\phi_{1}}$: additional benefit the decision-maker gains due to the successful improvement of individuals (as the successful improvement causes label change).
@@
    Loss from the failed improvement ${\phi_{2}}$: additional loss the decision-maker suffers due to the individuals’ failure to improve; this occurs because individuals who fail to improve only experience feature distribution shifts from $P_{X|Y}(x|0)$ to $P^{I}(x)$ but labels remain.
@@
    Loss from the manipulation $\phi_{3}$: additional loss the decision-maker suffers due to the successful manipulation of individuals; this occurs because individuals who manipulate successfully only change $P_{X|Y}(x|0)$ to $P_{X|Y}(x|1)$ but the labels remain unqualified.
@@ 3.1 Strategic decision-maker with adjusted preferences
    Original strategic decision-maker: the one with $k_{1}=k_{2}=k_{3}=u(1-\alpha)$ whose learning objective function $U$ follows (2.2); it considers both improvement and manipulation.
@@
The above examples show that a decision-maker, by changing the weights $k_{1},k_{2},k_{3}$ could find a policy that encourages certain types of individual behavior (as compared to the original policy $\theta^{*}$). Although the decision-maker can impact an individual’s behavior by adjusting its preferences via $k_{1},k_{2},k_{3}$, we emphasize that the actual utility it receives from the strategic individuals is always determined by $U(\theta)$ given in (2.2). Indeed, we can regard the framework with adjusted weights ((5)) as a regularization method. We discuss this in more detail in App. B.4.
@@ 4 Impacts of Adjusting Preferences
Next, we investigate the impacts of adjusting preferences. We aim to understand how a decision-maker by adjusting preferences (i.e., changing $k_{1},k_{2},k_{3}$) could affect the optimal policy (Sec. 4.1) and its fairness property (Sec. 4.3), as well as the resulting individual’s best response (Sec. 4.2).
@@ 4.1 Preferences shift the optimal threshold
We will start with the original strategic decision-maker (with $k_{1}=k_{2}=k_{3}=u(1-\alpha)$) whose objective function follows (2.2), and then investigate how adjusting preferences could affect the decision-maker’s optimal policy.
Complex nature of original strategic decision-maker. Unlike the non-strategic optimal policy, the analytical solution of strategic optimal policy that maximizes (2.2) is not easy to find. In fact, the utility function $U(\theta)$ of the original strategic decision-maker is highly complex, and the optimal strategic threshold $\theta^{*}$ may change significantly as $\alpha,F_{X|Y},F^{I},C_{M},C_{I},\epsilon,q$ vary. In App. C.2, we demonstrate the complexity of $U(\theta)$, which may change drastically as $\alpha,\epsilon,q$ vary. Although we cannot find the strategic optimal threshold precisely, we may still explore the impacts of decision-maker’s anticipation of strategic behavior on its policy (by comparing the strategic threshold $\theta^{*}$ with the non-strategic $\widehat{\theta}^{*}$), as stated in Thm. 4.1 below.
@@
Thm. 4.1 identifies a condition under which the strategic policy over-accepts individuals compared to the non-strategic one. Specifically, $\min_{\theta}P_{M}(\theta)\leq 0.5$ ensures that there exist policies under which the majority of individuals prefer improvement over manipulation. Intuitively, under this condition, strategic decision-maker by lowering the threshold (from $\widehat{\theta}^{*}$) may encourage more individuals to improve. Because $q$ is sufficiently large, more improvement brings more benefit to the decision-maker.
Optimal threshold under adjusted preferences. Despite the intricate nature of $U(\theta)$, the optimal strategic threshold may be shifted by adjusting the decision-maker’s preferences, i.e. changing the weights $k_{1},k_{2},k_{3}$ assigned to $\phi_{1},\phi_{2},\phi_{3}$ in (5). Next, we examine how the optimal threshold can be affected compared to the original strategic threshold by adjusting the decision-maker’s preferences. Denote $\theta^{*}(k_{i})$ as the strategic optimal threshold attained by adjusting weight $k_{i},i\in\{1,2,3\}$ of the original objective function $U(\theta)$. The results are summarized in Table 1. Specifically, the threshold gets lower as $k_{1}$ increase (Prop. 4.2). Adjusting $k_{2}$ or $k_{3}$ may result in the optimal threshold moving toward both directions, but we can identify sufficient conditions when adjusting $k_{2}$ or $k_{3}$ pushes the optimal threshold to move toward one direction (Prop. 4.3 and 4.4).

### v2

@@ Abstract
Machine learning systems have been widely used to make decisions about individuals who may behave strategically to receive favorable outcomes, e.g., they may genuinely improve the true labels or manipulate observable features directly to game the system without changing labels. Although both behaviors have been studied (often as two separate problems) in the literature, most works assume individuals can (i) perfectly foresee the outcomes of their behaviors when they best respond; (ii) change their features arbitrarily as long as it is affordable, and the costs they need to pay are deterministic functions of feature changes. In this paper, we consider a different setting and focus on imitative strategic behaviors with unforeseeable outcomes, i.e., individuals manipulate/improve by imitating the features of those with positive labels, but the induced feature changes are unforeseeable. We first propose a Stackelberg game to model the interplay between individuals and the decision-maker, under which we examine how the decision-maker’s ability to anticipate individual behavior affects its objective function and the individual’s best response. We show that the objective difference between the two can be decomposed into three interpretable terms, with each representing the decision-maker’s preference for a certain behavior. By exploring the roles of each term, we theoretically illustrate how a decision-maker with adjusted preferences may simultaneously disincentivize manipulation, incentivize improvement, and promote fairness. Such theoretical results provide a guideline for decision-makers to inform better and socially responsible decisions in practice.
@@ 1 Introduction
Individuals subject to algorithmic decisions often adapt their behaviors strategically to the decision rule to receive a desirable outcome. As machine learning is increasingly used to make decisions about humans, there has been a growing interest to develop learning methods that explicitly consider the strategic behavior of human agents. A line of research known as strategic classification studies this problem, in which individuals can modify their features at costs to receive favorable predictions. Depending on whether such feature changes are to improve the actual labels genuinely (i.e., improvement) or to game the algorithms maliciously (i.e., manipulation), existing works have largely focused on learning classifiers robust against manipulation (Hardt et al., 2016a) or designing incentive mechanisms to encourage improvement (Kleinberg & Raghavan, 2020; Bechavod et al., 2022; Xie et al., 2024). A few studies (Miller et al., 2020; Shavit et al., 2020; Horowitz & Rosenfeld, 2023) also consider the presence of both manipulation and improvement, where they exploit the causal structures of features and use structural causal models to capture the impacts of feature changes on labels.
To model the interplay between individuals and decision-maker, most existing works adopt (or extend based on) a Stackelberg game proposed by Hardt et al. (2016a), i.e., the decision-maker publishes its policy, following which individuals best respond to determine the modified features. However, these models (implicitly) rely on the following two assumptions that could make them unsuitable for certain applications: (i) individuals can perfectly foresee the outcomes of their behaviors when they best respond; (ii) individuals can change their features arbitrarily at costs, which are modeled as deterministic functions of the feature.
@@
Consider an example of college admission, where the students’ exam scores are treated as features in admission decisions. To get admitted, students may increase their scores by either cheating on exams (manipulation) or working hard (improvement). Here (i) individuals do not know the exact values of their original features (unrealized scores) and the modified features (actual scores received in an exam) when they best respond, but they have a good idea of what those score distributions would be like from their past experience; (ii) the cost of manipulation/improvement is not a function of feature change (e.g., students may cheat by hiring an imposter to take the exam and the cost of such behavior is more or less fixed). As the original feature was never realized, we cannot compute the feature change precisely and measure the cost based on it. Therefore, the existing models do not fit for these applications.
@@
Motivated by the above (more examples are also given in App. B.2), this paper studies strategic classification with unforeseeable outcomes. We first propose a novel Stackelberg game to model the interactions between individuals and the decision-maker. Compared to most existing models (Jagadeesan et al., 2021; Levanon & Rosenfeld, 2022), ours is a probabilistic framework that models the outcomes and costs of strategic behavior as random variables. Indeed, this framework is inspired by the models proposed in Zhang et al. (2022); Liu et al. (2020), which only considered either strategic manipulation (Zhang et al., 2022) or improvement (Liu et al., 2020). In contrast, our model significantly extends their works by considering both manipulation and improvement behaviors, investigating agents’ choices between them, and providing theoretical results and practical guideline for socially responsible decision-making in the new setting. Importantly, we focus on imitative strategic behavior where individuals manipulate/improve by imitating the features of those with positive labels, due to the following:
@@
  • It is inspired by imitative learning behavior in social learning, whereby new behaviors are acquired by copying social models’ behavior. It has been well-supported by literature in psychology and social science (Bandura, 1962; 1978). Recent works (Heidari et al., 2019; Raab & Liu, 2021) in ML also model individuals’ behaviors as imitating/replicating the profiles of their social models to study the impacts of fairness interventions.
@@
  • Decision-makers can detect easy-to-manipulate features (Bechavod et al., 2021) and stop using them when making decisions, so individuals can barely manipulate their features by themselves without changing labels. A better option for them is to mimic others’ profiles. Such imitation-based manipulative behavior is very common in the real world (e.g., cheating, identity theft) and even becomes increasingly worrying during recent years [^1] As COVID-19 hit the world, candidates are more commonly permitted to take exams/assessments (e.g., GRE, TOEFL, or online assessments of companies) remotely. Although many institutions are diligent in designing novel challenges to prevent candidates from directly finding the answers on the internet, the remote nature makes it easier to hire qualified imposters to take the assessments instead of them. Talha (2024) illustrated how students can let others take the GRE instead of them when the test is permitted to be taken at home..
@@
    We study how adjusting the decision-maker’s preferences can affect the optimal policy and its fairness property, as well as the resulting individual’s best response (Sec. 4). We also illustrate how the decision-maker can adjust preferences to disincentivize manipulation, incentivize improvement and promote fairness in practice (Sec. 4 and App. B.5).
@@ 2 Problem Formulation
Consider a group of individuals subject to some ML decisions. Each individual has an observable feature $X\in\mathbb{R}$ and a hidden label $Y\in\{0,1\}$ indicating its qualification state (“0" being unqualified and “1" being qualified). [^2] Similar to prior work (Zhang et al., 2022; Liu et al., 2018), we present our model in one-dimensional feature space. Note that our model and results are applicable to high dimensional space, in which individuals imitate and change all features as a whole based on the joint conditional distribution $P_{X|Y}$ regardless of the dimension of $X$. The costs can be regarded as the sum of an individual’s effort to change features in all dimensions. Let $\alpha:=\Pr(Y=1)$ be the population’s qualification rate, and $P_{X|Y}(x|1)$, $P_{X|Y}(x|0)$ be the feature distributions of qualified and unqualified individuals, respectively. A decision-maker makes binary decisions $D\in\{0,1\}$ (“0" being reject and “1" being accept) about individuals based on a threshold policy with acceptance threshold $\theta\in\mathbb{R}$: $\pi(x)=P_{D|X}(1|x)=\textbf{1}(x\geq\theta)$. To receive positive decisions, individuals with information of policy $\pi$ may behave strategically by either manipulating their features or improving the actual qualifications. [^3] We assume all unqualified individuals who stay in the decision-making system either manipulate or improve. In other words, they already have a sufficiently large initial utility by staying in the system and trying to be qualified. We discuss the details in App. B.3. Formally, let $M\in\{0,1\}$ denote individual’s action, with $M=1$ being manipulation and $M=0$ being improvement.
Outcomes of strategic behavior. Both manipulation and improvement result in the shifts of feature distribution. Specifically, for individuals who choose to manipulate, we assume they manipulate by “stealing" the features of those qualified (Zhang et al., 2022), e.g., students cheat on exams by hiring qualified imposters. Moreover, we assume the decision-maker can identify the manipulation behavior with probability $\epsilon\in[0,1]$ (Estornell et al., 2021). Individuals, once getting caught manipulating, will be rejected directly. For those who decide to improve, they work hard to imitate the features of those qualified (Bandura, 1962; Raab & Liu, 2021; Heidari et al., 2019). With probability $q\in[0,1]$, they improve the label successfully (overall $\alpha$ increases) and the features conform the distribution $P_{X|Y}(x|1)$; with probability $1-q$, they slightly improve the features but fail to change the labels, and the improved features conform a new distribution $P^{I}(x)$. Throughout the paper, we make the following assumption on feature distributions.
@@
Assumption 2.1 is relatively mild and has been widely used (e.g., (Tsirtsis et al., 2019; Zhang et al., 2020b)). It can be satisfied by a wide range of distributions (e.g., exponential, Gaussian) and the real data (e.g., FICO data used in Sec. 5). It implies that an individual is more likely to be qualified as feature value increases. Meanwhile, compared to the unqualified individuals, the individuals who improve but fail also tend to have higher feature values. Individuals have a good knowledge of their true qualifications by observing their peers or previous individuals who received positive decisions Raab & Liu (2021), and only unqualified individuals have incentives to take action Dong et al. (2018) since $P_{X|Y}(x|1)$ is always the best attainable outcome (as manipulation and improvement only bring additional cost but no benefit to qualified individuals).
@@
The above formulation captures the imitative strategic behavior with unforeseeable outcomes (e.g., college admission example in Sec. 1): individuals best respond based on feature distributions but not the realizations, and the imitation costs (e.g., hiring an imposter) for individuals from the same group follow the same distribution (Liu et al., 2020), as opposed to being a function of feature changes.  equation 1 above can further be written based on CDF of $C_{M}-C_{I}$, i.e., the difference between manipulation and improvement costs. We make the following assumption on its PDF.
@@
Under Assumption 2.2, $P_{M}(\theta)$ is continuous and satisfies the following: (i) If $q+\epsilon\geq 1$, then $P_{M}(\theta)$ strictly increases. (ii) If $q+\epsilon<1$, then $P_{M}(\theta)$ first increases and then decreases, thereby existing a unique maximizer $\theta_{max}$. Moreover, maximizer $\theta_{max}$ increases in $q$ and $\epsilon$.
@@ 2.2 Decision-maker’s optimal policy
Strategic optimal policy. Given cost and feature distributions, a strategic decision-maker can anticipate an individual’s best response (equation 1) and incorporate it in determining its optimal policy. Under a threshold policy $\pi(x)=\textbf{1}(x\geq\theta)$, the objective $U(\pi)$ can be written as a function of $\theta$, i.e.,
@@ 3 Decomposition of the Objective Difference
As shown in equation 4, the objective difference $\Phi$ can be decomposed into three terms $\phi_{1},\phi_{2},\phi_{3}$. It turns out that each term is interpretable and indicates the impact of a certain type of individual behavior on the decision-maker’s utility. We discuss these in detail as follows.
@@
    Benefit from the successful improvement ${\phi_{1}}$: additional benefit the decision-maker gains due to the successful improvement of individuals (as the successful improvement causes label change). In the college admission example, $\phi_{1}$ stands for the utility gain caused by students studying hard to be qualified.
@@
    Loss from the failed improvement ${\phi_{2}}$: additional loss the decision-maker suffers due to the individuals’ failure to improve; this occurs because individuals who fail to improve only experience feature distribution shifts from $P_{X|Y}(x|0)$ to $P^{I}(x)$ but labels remain. In the college admission example, $\phi_{2}$ stands for the utility loss caused by students who seem to try but fail to be qualified.
@@
    Loss from the manipulation $\phi_{3}$: additional loss the decision-maker suffers due to the successful manipulation of individuals; this occurs because individuals who manipulate successfully only change $P_{X|Y}(x|0)$ to $P_{X|Y}(x|1)$ but the labels remain unqualified. In the college admission example, $\phi_{3}$ stands for the utility loss caused by students who cheat/hire imposters.
@@ 3.1 Strategic decision-maker with adjusted preferences
    Original strategic decision-maker: the one with $k_{1}=k_{2}=k_{3}=u(1-\alpha)$ whose learning objective function $U$ follows equation 2.2; it considers both improvement and manipulation.
@@
The above examples show that a decision-maker, by changing the weights $k_{1},k_{2},k_{3}$ could find a policy that encourages certain types of individual behavior (as compared to the original policy $\theta^{*}$). Although the decision-maker can impact an individual’s behavior by adjusting its preferences via $k_{1},k_{2},k_{3}$, we emphasize that the actual utility it receives from the strategic individuals is always determined by $U(\theta)$ given in equation 2.2. Indeed, we can regard the framework with adjusted weights (equation 5) as a regularization method. We discuss this in more detail in App. B.4.
@@ 4 Impacts of Adjusting Preferences
Next, we investigate the impacts of adjusting preferences. We aim to understand how a decision-maker can adjust preferences (i.e., changing $k_{1},k_{2},k_{3}$) to affect the optimal policy (Sec. 4.1) and its fairness property (Sec. 4.3), as well as the resulting individual’s best response (Sec. 4.2).
@@ 4.1 Preferences shift the optimal threshold
We will start with the original strategic decision-maker (with $k_{1}=k_{2}=k_{3}=u(1-\alpha)$) whose objective function follows equation 2.2, and then investigate how adjusting preferences could affect the decision-maker’s optimal policy.
Complex nature of original strategic decision-maker. Unlike the non-strategic optimal policy, the analytical solution of strategic optimal policy that maximizes equation 2.2 is not easy to find. In fact, the utility function $U(\theta)$ of the original strategic decision-maker is highly complex, and the optimal strategic threshold $\theta^{*}$ may change significantly as $\alpha,F_{X|Y},F^{I},C_{M},C_{I},\epsilon,q$ vary. In App. C.2, we demonstrate the complexity of $U(\theta)$, which may change drastically as $\alpha,\epsilon,q$ vary. Although we cannot find the strategic optimal threshold precisely, we may still explore the impacts of the decision-maker’s anticipation of strategic behavior on its policy (by comparing the strategic threshold $\theta^{*}$ with the non-strategic $\widehat{\theta}^{*}$), as stated in Thm. 4.1 below.
@@
Thm. 4.1 identifies a condition under which the strategic policy over-accepts individuals compared to the non-strategic one. Specifically, $\min_{\theta}P_{M}(\theta)\leq 0.5$ ensures that there exist policies under which the majority of individuals prefer improvement over manipulation. Intuitively, under this condition, a strategic decision-maker by lowering the threshold (from $\widehat{\theta}^{*}$) may encourage more individuals to improve. Because $q$ is sufficiently large, more improvement brings more benefit to the decision-maker.
Optimal threshold under adjusted preferences. Despite the intricate nature of $U(\theta)$, the optimal strategic threshold may be shifted by adjusting the decision-maker’s preferences, i.e. changing the weights $k_{1},k_{2},k_{3}$ assigned to $\phi_{1},\phi_{2},\phi_{3}$ in equation 5. Next, we examine how the optimal threshold can be affected compared to the original strategic threshold by adjusting the decision-maker’s preferences. Denote $\theta^{*}(k_{i})$ as the strategic optimal threshold attained by adjusting weight $k_{i},i\in\{1,2,3\}$ of the original objective function $U(\theta)$. The results are summarized in Table 1. Specifically, the threshold gets lower as $k_{1}$ increases (Prop. 4.2). Adjusting $k_{2}$ or $k_{3}$ may result in the optimal threshold moving toward both directions, but we can identify sufficient conditions when adjusting $k_{2}$ or $k_{3}$ pushes the optimal threshold to move toward one direction (Prop. 4.3 and 4.4).